140804 LML Observations

Isaac arrived at LML at 9:30am. The doors to the collector had deployed and a tiny sample of approximately 5ml was visible in the collection jar. Since the sample was of minimal volume it was discarded. The collector was then sprayed with 409 and rinsed with DI water. the collector was additionally disassembled and its components cleaned with 409 and DI water. The collector was then reassembled for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *